The problem is: An alloy of metals is 25% copper. Another alloy is 50% copper. How much should be used to make 1000 grams of an alloy that is 45% copper? I am having a problem because in the problem it doesn't state how much is in the first two alloys.
Answer by ankor@dixie-net.com(22740)

The problem is: An alloy of metals is 25% copper.
Another alloy is 50% copper.
How much should be used to make 1000 grams of an alloy that is 45% copper?
:
Let x = amt of 50% copper to be used
then resulting mixture is to be 1000 gr, therefore:
(1000-x) = amt of 25% copper required
:
Write a mixture equation using the decimal equiv
.50x + .25(1000-x) = .45(1000)
.50x + 250 - .25x = 450
.50x - .25x = 450 - 250
.25x = 200
x = 200/.25
x = 800 grams of 50% copper used
:
I'll let you find the amt of 25% copper is required